OpenStudy (anonymous):

Give the equivalent numerator if the denominator is (a - 6)(a - 1)(a + 6).

OpenStudy (anonymous):

\[6/a ^{2}-7a+6\]

myininaya (myininaya):

\[\frac{?}{(a-6)(a-1)(a+6)}\]=whatever fraction you have then just muliply both sides by (a-6)(a-1)(a+6) to find what ? is

OpenStudy (anonymous):

you multiply (a - 6)(a - 1)(a + 6) by 6

OpenStudy (anonymous):

but i dont know what that is

myininaya (myininaya):

6(a-6)(a-1)(a+6)

OpenStudy (anonymous):

yeah but i dont know the answer

myininaya (myininaya):

what is the fraction 6/a^2-7a+6 or 6/(a^2-7a+6)

myininaya (myininaya):

\[\frac{6}{a^2-7a+6}=\frac{?}{(a-1)(a-6)(a+6)}\] \[\frac{6}{(a-6)(a-1)}=\frac{?}{(a-1)(a-6)(a+6)}\] now what does the ? need to be for both sides to be the same? ?=6(a+6) right?

OpenStudy (anonymous):

yeah maybe?? hah im so confused

OpenStudy (anonymous):

yes that was right!!!